OpenStudy (anonymous):

How many grams of beryllium chloride are needed to make 125 mL of a 0.050 M solution?

OpenStudy (chmvijay):

beryllium chloride molar mass can u tell me ?

OpenStudy (anonymous):

molecular mass of beryllium chloride = (9)+(35.5*2)=9+71=80 gm 0.050 M Solution means in 1000 ml 0.05 moles dissolved. 1000ml has 0.05 Moles of BeCl2= 0.05*80=4gm 125 gm has (4*125)/1000=0.5gm(ans)